7 years ago, my sister was twice the age I was then plus two years making her 32
The combined age of us two years ago was 57 How old am I now ?

Answer by Boreal(15235)

You are x
sister is y
x-2+y-2=57
x+y=61
2(x-7)+2=32
2x-14=30
2x=44
x=22 your age ANSWER
y=39, sister's age
2 years ago they were 20 and 37=57
7 years ago, sister was 32 and you were 15--32 is 2 x 15 +2.